OpenStudy (yrelhan4):

Four fair dice D1, D2, D3and D4 each having six faces numbered 1,2,3,4,5 and 6 are rolled simultaneously. The probability that D4 shows a number appearing on one of D1 , D2and D3 is

OpenStudy (yrelhan4):

@shubhamsrg @Mashy

OpenStudy (kropot72):

The chance that D4 shows a number that appears on only one of D1, D2 and D3 is the chance that the number appears on one times the chance that the number does not appear on the other two dice: \[P(only\ one)=\frac{1}{6}\times \frac{5}{6}\times \frac{5}{6}\]

OpenStudy (yrelhan4):

P(two)=1/6*1/6*5/6 follow this for p(3) and add?

OpenStudy (yrelhan4):

@kropot72

OpenStudy (yrelhan4):

i got it. 1-(5/6)^3 would give me the answer. thank you for replying. :)